OpenStudy (olivya999):

Which of the following represents fifth root of x to the fourth power in exponential form?

OpenStudy (olivya999):

\[\sqrt[5]{x ^{4}}\]

OpenStudy (olivya999):

A) 4x^5 b) 5x^4 C) \[x ^{\frac{ 5 }{ 4 }}\] D) \[x ^{\frac{ 4 }{ 5}}\]

OpenStudy (latinc):

same as (x^4)^(1/5) or x^(4/5)

OpenStudy (latinc):

5th is your root so 5 would be at the bottom

OpenStudy (olivya999):

So D?

OpenStudy (latinc):

YEs Sir:)

OpenStudy (olivya999):

Thank you